OpenStudy (anonymous):

What is the greatest common factor of 90xy3 and 12xy?

OpenStudy (anonymous):

90 = 2* 3^2 * 5 12 = 2^2 *3 gcf = 6 so gcf is 6xy just choose the factors they have in common with the least exp.

OpenStudy (anonymous):

dpalnc is right, the gcf is 6xy.

jhonyy9 (jhonyy9):

yes 6xy --- is sure right